# Process-based Inference for Spatial Energetics Using Bayesian Predictive Stacking

### v2

@@ 1. Introduction
Spatial energetics is a rapidly emerging area in biomedical and health sciences that aims to examine how environmental characteristics, space, and time are linked to activity-related health behaviors (James et al., 2016). Examples include, but are not limited to, using data from wearable devices as biomarkers and risk factors in studying adverse health outcomes for respiratory health. Inferential objectives for spatial energetics comprise two exercises: (i) estimate measured health variables, typically related to metabolic activities, over paths or trajectories traversed by subjects as they conduct their daily physical activities; and (ii) predict the health variables for a subject at arbitrary trajectories. Spatial-temporal process models seem a natural choice as they use space-time coordinates from Global Positioning Systems (GPS) embedded within actigraph units.
@@
Some salient features of spatial energetics require consideration. Unlike in some clinical studies associated with mobile health data where only the temporal nature of streaming data is of inferential interest, here inferential interest centers around estimation and prediction of metabolic measurements over arbitrary spatial trajectories or paths. This differs from customary geostatistics and spatial-temporal data analysis (see, e.g., Banerjee et al., 2014; Cressie and Wikle, 2015, and references therein) where statistical inference proceeds from spatial-temporal processes $\{w(s,t):s\in\mathcal{S},\ t\in\mathcal{T}\}$, where $\mathcal{S}\subset\mathbb{R}^{d}$ with $d=2$ or $3$ and $\mathcal{T}\subset\mathbb{R}^{+}\cup\{0\}$. For mobile health applications, the spatial domain is typically an arbitrary string of spatial coordinates defining the path or trajectory traversed by the subject. This trajectory is completely arbitrary and need not enjoy mathematically attractive features as are available for Riemannian manifolds to carry out inference (Li et al., 2023). Furthermore, treated as continuously evolving over time, the spatial coordinates are best considered as functions of time.
The rapidly emerging literature on statistical analysis of streaming wearable device data has almost exclusively focused on longitudinal models and purely temporal processes (Chang and McKeague, 2022; Luo et al., 2023; Banker and Song, 2023). Their analytical objectives are not concerned with the spatial attributes of trajectories. Given the nature of streaming spatial-temporal data, we explore models that treat (i) space as continuous and time as discrete, and (ii) both space and time as continuous. The terms “discrete” and “continuous” describe whether inference is sought at the same scale where the data are available or at arbitrary resolutions. The former is usually analyzed using spatial-temporal dynamic models (see, e.g., Stroud et al., 2001; Peña and Poncela, 2004; Gamerman et al., 2008), while the latter employs spatial-temporal processes specified using appropriate covariance kernels. Gaussian processes are conspicuous in spatial-temporal data analysis, but have largely focused on Euclidean domains and, more generally, on compact Riemannian manifolds (see, e.g., Li et al., 2023). Mobility data, on the other hand, arise from completely arbitrary trajectories that do not satisfy the conditions on manifolds. Related literature on non-Euclidean domains includes Hoef et al. (2006), Hoef and Peterson (2010) and Santos-Fernandez et al. (2022), who considered spatial modeling of data from rivers and streams, Hooten et al. (2017), who reviewed spatial (discrete & continuous) temporal modeling of the trajectory of animal movement, and Jurek et al. (2023), who analyzed mobility trajectories under flight-pause models.
Our study differs from the aforementioned studies. While flight-pause models and animal movement models seek inference on the evolution of the path itself, we seek to analyze data on variables of interest that have been collected at high resolutions on subjects moving along trajectories. Second, unlike the modeling of phenomena on fixed geographic structures, such as river networks, where the domain of interest is modeled effectively as a fixed graph, our inferential interest lies in predicting the variables over arbitrary trajectories over whose shape or structure we have no control. We seek fully model-based uncertainty quantification in our inference, which will include predicting variables on hypothetical paths that have not generated any data yet.
We underscore the inferential difficulties inherent in such models. Most notably, the stochastic process parameters are not identifiable and are difficult to estimate from finite realizations of the process (Zhang, 2004; Tang et al., 2021). This is manifested by poor convergence of samples from iterative Markov chain Monte Carlo (MCMC) algorithms or numerical instabilities in other iterative algorithms such as Integrated Nested Laplace Approximations or Variational Bayes (see, e.g., Robert and Casella, 2004; Rue et al., 2009; Gelman et al., 2013; Murphy, 2023, for details on iterative Bayesian computational algorithms). Therefore, we develop and execute a computationally efficient Bayesian stacking approach for fast inference that relies on fixing some parameters to achieve analytically tractable distribution theory and then “stack” over these analytical posterior distributions to obtain an averaged posterior (Yao et al., 2018; Bhatt et al., 2017; Zhang et al., 2023).
@@
The balance of our paper proceeds as follows. Section 2 provides an overview of Bayesian hierarchical models that treat space as continuous and time as discrete. In particular, we show how to adapt familiar Bayesian dynamic linear models (West and Harrison, 1997; Prado et al., 2021; Stroud et al., 2001) to actigraph data. Sections 3 develops a spatial-temporal process model (Cressie and Huang, 1999; Stein, 2005; Gneiting, 2002) that treats space and time as continuous. Section 4 develops predictive stacking algorithms for the models we develop by averaging over sets of conjugate Bayesian models with accessible posterior distributions. Section 5 collects some results on distribution theory and offers some theoretical insights. Sections 6 and 7 present simulation experiments and an illustrative application, respectively, for our methods. Section 8 concludes the article with a discussion and pointers to future research.
@@ 2. Continuous space and discrete time models
Broadly speaking, spatial-temporal models are classified according to whether space and time are modeled as continuous or discrete processes. Bayesian dynamic linear models, or DLMs (West and Harrison, 1997; Prado et al., 2021), are widely employed for analyzing temporal data by modeling time over a countable set of integers and space using a continuous random field evolving over the time steps. We first show how such models can be effectively employed for actigraph data.
@@ 2.1. Spatial-temporal Bayesian DLMs
Two adaptations are relevant for spatial energetics. First, $y_{t}$ consists of $n_{t}$ measurements recorded at epoch $t$ independently over a group of subjects. The collected measurements $\{y_{t}:t\in\mathcal{T}\}$ is called an actigraph time-sheet, where each epoch also provides values for the elements of $X_{t}$. Our available data, therefore, is $\{(y_{t},X_{t}):t\in\mathcal{T}\}$. While each epoch implicitly contains information on the spatial locations for the subjects’ measurements, the inferential goals for these population-level studies do not entail spatial attributes and, instead, are concerned with inferring about relationships between metabolic measurements (representing levels of physical activity) and environmental variables (representing green spaces, climate and weather, local topography, nature of activity being performed by the subject, etc.). It is reasonable to assume that $V_{t}$ is diagonal since measurements on subjects are taken independently of each other and shared features across subjects are accounted for with explanatory variables and random effects in $F_{t}$. The covariance matrix for the elements of $\theta_{t}$, $S_{t}$, is assumed to be diagonal if latent associations are adequately accounted for by $G_{t}$. Alternative models could specify $S_{t}$ from design considerations or be modeled using an appropriate prior distribution (see, e.g., West and Harrison, 1997; Prado et al., 2021).
@@
The model in (1) readily facilitates comprehensive inference through MCMC or forward filtering-backward sampling (Carter and Kohn, 1994; Frühwirth-Schnatter, 1994). However, with high-dimensional parameters, these methods require substantial computational resources that render their practical application to be challenging or even infeasible. Consequently, we devise Bayesian predictive stacking that exploits analytically tractable posterior distributions.
@@ 2.2. Dynamic trajectory model
$y_{t}(\gamma(t))={x}_{t}(\gamma(t))^{\mathrm{\scriptscriptstyle T}}{\beta}_{t}+w_{t}(\gamma(t))+\eta_{1t}(\gamma(t)),\quad\eta_{1t}(\gamma(t))\overset{\mathrm{i.i.d.}}{\sim}N(0,\sigma^{2})\quad\mbox{for}\quad t=1,\ldots,T,$ (2)
where ${x}_{t}(\gamma(t))$ is a $p$-dimensional explanatory variable, ${\beta}_{t}$ is a $p$-dimensional time-varying regression coefficient, $w_{t}(\gamma(t))$ and $\eta_{1t}$ are zero-centered spatial and white noise processes, respectively, at time $t$, and $\sigma^{2}$ is the variance of the white noise process. The domain of the process $w_{t}(\cdot)$ is not Euclidean, but an arbitrary trajectory defined by a string of coordinates mapped by $\gamma(t)$. Furthermore, the subject may revisit the same location a number of times, yielding multiple values at the same location $\gamma(t)$, making the spatial covariance matrix singular and, therefore, precluding legitimate probabilistic inference. We obviate this as follows.
Let $\Gamma=\{\gamma(1),\ldots,\gamma(T)\}$ be a complete enumeration of spatial locations visited by the subject, of which $n\leq T$ are distinct spatial locations, and let $\widetilde{\Gamma}=\{\widetilde{\gamma}_{1},\ldots,\widetilde{\gamma}_{n}\}\subseteq\Gamma$ be the subset of distinct locations. We define a latent process $z_{t}(\gamma(t))$ and map $w_{t}(\gamma(t))=\sum_{j=1}^{n}b(\gamma(t),\widetilde{\gamma}_{j})z_{t}(\widetilde{\gamma}_{j})$, where $b(\gamma(t),\widetilde{\gamma}_{j}):\Gamma\times\widetilde{\Gamma}\to\{0,1\}$ such that $b(\gamma(t),\widetilde{\gamma}_{j})=1$ if $\gamma(t)=\widetilde{\gamma}_{j}$ and 00 otherwise. This yields $w=Bz$, where $w=(w_{1}(\gamma(1)),\ldots,w_{T}(\gamma(T)))^{\mathrm{\scriptscriptstyle T}}$ is $T\times 1$, ${z}=({z}_{1}^{\mathrm{\scriptscriptstyle T}},\ldots,{z}_{T}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ is $nT\times 1$ with each ${z}_{t}=({z}_{t}(\widetilde{\gamma}_{1}),\ldots,{z}_{t}(\widetilde{\gamma}_{n}))^{\mathrm{\scriptscriptstyle T}}$ being $n\times 1$, and $B$ is $T\times nT$, whose $(t,n(t-1)+j)$th entry is $b(\gamma(t),\widetilde{\gamma}_{j})$. This formulates a map of the latent spatial effects across whole spaces and times, $z$, to those at the observed points, $w$. If ${\beta}=({\beta}_{1}^{\mathrm{\scriptscriptstyle T}},\ldots,{\beta}_{T}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$, then temporal autoregressive models for $\beta$ and $z$ are specified as ${\beta}=(A\otimes I_{p}){\beta}+\eta_{2t}$ and $z=(A\otimes I_{n})z+\eta_{3t}$, respectively, where $\eta_{2t}\overset{\mathrm{i.i.d.}}{\sim}N({0},\sigma^{2}\delta_{\beta}^{2}(I_{T}\otimes W_{p}))$ and $\eta_{3t}\sim N(0,\sigma^{2}\delta_{z}^{2}(I_{T}\otimes K_{\phi}))$, $A=\left(\begin{smallmatrix}{0}^{\mathrm{\scriptscriptstyle T}}&0\\ I_{T-1}&0\\ \end{smallmatrix}\right)\in\mathbb{R}^{T\times T}$, $W_{p}\in\mathbb{R}^{p\times p}$ is a correlation matrix among the coefficients and $\otimes$ denotes the Kronecker product. We construct the following augmented model,
@@
3. Continuous space and continuous time trajectory model
We can treat actigraph data as a partial realization of a continuous spatial-temporal process. We write $y(\gamma(t),t)$ to be the measurement that can exist, at a conceptual level, at any time $t\in\mathbb{R}^{+}$ and on a continuous geographic point $\gamma(t)\in\mathbb{R}^{2}$ at $t$. We define the following regression model over space-time coordinates $(\gamma(t),t)$ generated by a finite collection of $n$ time points,
@@ 3. Continuous space and continuous time trajectory model
where ${x}(\gamma(t),t)$ is $p\times 1$ consisting of explanatory variables, ${\beta}(t)$ is the corresponding $p\times 1$ vector of slopes, $z(\gamma(t),t)$ is a zero-centered spatial-temporal process and $\eta_{1}(t)$ is the measurement error distributed as a zero-centered Gaussian distribution with variance $\sigma^{2}$.
@@
where the correlation kernel is
@@
which represents the spatial-temporal correlation of the data. For each regression coefficient, we consider the following process:
@@
with the temporal correlation kernel $C_{\xi}(t,t^{\prime})=\exp(-\xi^{2}|t-t^{\prime}|^{2})$. We note that $K_{\phi}$ is a positive-definite kernel, ensuring that the stochastic process (5) is well-defined. In particular, it is crucial to note that even when $\gamma(t)=\gamma(t^{\prime})$, i.e., the subject returns to the same location at a later time, the function $(1+\phi_{1}|t-t^{\prime}|^{2})^{-1}$ is positive-definite. This is seen by noting that
@@
Because $|t-t^{\prime}|^{2}$ and $1$ are conditionally negative-definite, the integrand is positive-definite from Schoenberg’s theorem (e.g., Phillips et al., 2019), which implies that $(1+\phi_{1}|t-t^{\prime}|^{2})^{-1}$ is also positive-definite. Assume that we observe $y$ within finite space-time points $(\Gamma,\mathcal{T})=\{(\gamma(t_{i}),t_{i})\mid i=1,\ldots,n\}\subset\mathbb{R}^{2}\times\mathbb{R}^{+}$. Then, (4),(5) and (7) for $n$ space-time points results in the linear system:
@@
where each ${x}_{j}\in\mathbb{R}^{n\times n}$ is diagonal with entries $x_{j}(\gamma(t_{i}),(t_{i}))$ for $i=1,\ldots,n$ and $j=1,\ldots,p$; $X$ is $3n\times(p+1)n$, ${\theta}$ is $(p+1)n\times 1$ consisting of the $pn\times 1$ vector of regression coefficients $\beta=(\beta_{1}^{\mathrm{\scriptscriptstyle T}},\ldots,\beta_{p}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ and the $n\times 1$ vector $z=(z(\gamma(t_{1}),t_{1}),\ldots,z(\gamma(t_{n}),t_{n}))^{\mathrm{\scriptscriptstyle T}}$, $K_{\phi}(\Gamma,\mathcal{T})=(K_{\phi}((\gamma(t_{i}),t_{i}),(\gamma(t_{j}),t_{j})))$ and $C_{\xi}(\mathcal{T})=(C_{\xi}(t_{i},t_{j}))$ are both $n\times n$, where $i,j=1,\ldots,n$ and $S=I_{n}\oplus\left(I_{p}\otimes\delta_{\beta}^{2}C_{\xi}(\mathcal{T})\right)\oplus\delta_{z}^{2}K_{\phi}(\Gamma,\mathcal{T})$. We further assign the prior distribution $\sigma^{2}\sim IG(a_{\sigma},b_{\sigma})$, where $a_{\sigma}$ and $b_{\sigma}$ are fixed rate and scale parameters for the inverse-Gamma distribution. As in (3), the prior distribution for $\theta$ is absorbed into (8) and the posterior distribution for $\{\theta,\sigma^{2}\}$ for any fixed set $\{\delta_{\beta},\delta_{z},\xi,\phi\}$ is in the Normal-IG family. We exploit these familiar distributions to devise stacked inference for the processes $\beta_{j}(t)$ and $z(\gamma(t),t)$.

### v3

@@ 1. Introduction
Spatial energetics is a rapidly emerging area in biomedical and health sciences that aims to examine how environmental characteristics, space, and time are linked to activity-related health behaviors (James et al., 2016; Bürgi et al., 2016). Examples include, but are not limited to, using data from wearable devices as biomarkers and risk factors in studying adverse health outcomes for respiratory health (Kim et al., 2024). Inferential objectives for spatial energetics comprise two exercises: (i) estimate measured health variables, typically related to metabolic activities, over paths or trajectories traversed by subjects while conducting their daily physical activities; and (ii) predict health variables for a subject at arbitrary trajectories. Spatial-temporal process models seem a natural choice as they use space-time coordinates from Global Positioning Systems (GPS) embedded in actigraph units.
@@
Some salient features of spatial energetics require consideration. Unlike in some clinical studies associated with mobile health data, where only the temporal nature of streaming data is of inferential interest, here inferential interest centers around estimation and prediction of metabolic measurements over arbitrary spatial trajectories or paths. This differs from customary geostatistics and spatial-temporal data analysis (see, e.g., Banerjee et al., 2014; Cressie and Wikle, 2015, and references therein) where statistical inference proceeds from spatial-temporal processes $\{w(s,t):s\in\mathcal{S},\ t\in\mathcal{T}\}$, where $\mathcal{S}\subset\mathbb{R}^{d}$ with $d=2$ or $3$ and $\mathcal{T}\subset\mathbb{R}^{+}\cup\{0\}$. For mobile health applications, the spatial domain is typically an arbitrary string of spatial coordinates defining the path or trajectory traversed by the subject. This trajectory is completely arbitrary and need not enjoy mathematically attractive features as are available for Riemannian manifolds to carry out inference (Li et al., 2023). Importantly, treated as continuously evolving over time, the spatial coordinates are functions of time.
The emerging literature on actigraph data has largely focused on temporal modeling (Crouter et al., 2006a; Chang and McKeague, 2022; Luo et al., 2023; Banker and Song, 2023) and have excluded spatial attributes of trajectories. Spatial extensions seek stochastic processes using either dynamic models (Stroud et al., 2001; Peña and Poncela, 2004; Gamerman et al., 2008) or continuous processes using covariance kernels. Gaussian processes are conspicuous, but have largely focused on Euclidean domains. Actigraph data, however, arise from arbitrary trajectories that do not offer covariance kernels with exploitable geometric properties. While inference on non-Euclidean domains is possible (Ver Hoef et al., 2006; Ver Hoef and Peterson, 2010; Santos-Fernandez et al., 2022), they do not consider spatial coordinates as functions over time. This also preclude predicting metabolic variables at arbitrary space-time coordinates, which is central to our aims.
We proffer our specific contributions here. While related, our current developments differ from Hooten et al. (2017), who modeled spatial-temporal animal movement, and Jurek et al. (2023), who analyzed movement using flight-pause models. Such models seek inference on the evolving path itself, whereas we analyze data that are collected at high resolutions on subjects moving along trajectories. Unlike studies over fixed geographic structures, where the domain of interest is a fixed graph and one does not pursue inference beyond a fixed structure, we seek predictive inference on metabolic activity would subjects adopt posited new trajectories for their physical activity. We devise Bayesian inference, including predictions over trajectories, with fully model-based uncertainty quantification for mobile data. Rather than opting for excessively complex non-linear models and potentially cumbersome Markov chain Monte Carlo (MCMC) (with which we compare) or other iterative algorithms such as Integrated Nested Laplace Approximations or Variational Bayes (see, e.g., Robert and Casella, 2004; Rue et al., 2009; Gelman et al., 2013; Murphy, 2023) for fitting them, we extend predictive stacking (Yao et al., 2018; Bhatt et al., 2017; Zhang et al., 2023) for fast Bayesian inference on actigraph data from wearables.
Here, we overcome two specific challenges: (i) we avoid degenerate stochastic processes that would normally result from a subject visiting the same location multiple times (typical in mobile health data from walking, running); and (ii) accommodating spatial-temporal dependence accounting for locations being functions of time. Notably, we stack over analytically tractable posterior distributions to achieve fast and robust inference, which also renders our learning framework suitable for artificially intelligent “recommender” systems based upon an individual’s specific health attributes.
@@
Section 2 discusses Bayesian hierarchical models that treat space as continuous and time as discrete. In particular, we show how to adapt Bayesian dynamic linear models (West and Harrison, 1997; Prado et al., 2021; Stroud et al., 2001) to actigraph data. Sections 3 develops a spatial-temporal process model (Cressie and Huang, 1999; Stein, 2005; Gneiting, 2002) treating both space and time as continuous. Section 4 develops predictive stacking by averaging over sets of conjugate Bayesian models with accessible posterior distributions. Section 5 collects some results on distribution theory and offers theoretical insights. Sections 6 and 7 present simulation experiments and an illustrative application, respectively. Section 8 concludes the article with a discussion and pointers to future research.
@@ 2. Continuous space and discrete time models
Spatial-temporal models are classified based on continuous or discrete processes for space and time. Bayesian dynamic linear models (DLMs, West and Harrison, 1997; Prado et al., 2021) analyze temporal data using discrete time steps. Space is modeled as a continuous random field evolving over time. We first show how such models can be adapted to actigraph data after which we consider spatial-temporal covariance kernels for actigraph data.
@@ 2.1. Spatial-temporal Bayesian DLMs
Two adaptations are relevant for spatial energetics. First, $y_{t}$ consists of $n_{t}$ measurements recorded at epoch $t$ independently over a group of subjects. The collected measurements $\{y_{t}:t\in\mathcal{T}\}$ constitute an actigraph time-sheet, where each epoch provides values for the elements of $X_{t}$. Our available data, therefore, is $\{(y_{t},X_{t}):t\in\mathcal{T}\}$. While each epoch implicitly contains information on the spatial locations for the subjects’ measurements, the inferential goals for these population-level studies do not entail spatial attributes and, instead, are concerned with inferring about relationships between metabolic measurements (representing levels of physical activity) and environmental variables (representing green spaces, climate and weather, local topography, nature of activity being performed by the subject, etc.). We reasonably assume that $V_{t}$ is diagonal since measurements on subjects are independent and shared features across subjects are accounted for in explanatory variables and random effects in $F_{t}$. The covariance matrix for the elements of $\theta_{t}$, $S_{t}$, is assumed to be diagonal if latent associations are adequately accounted for by $G_{t}$. Alternative models could specify $S_{t}$ from design considerations or be modeled using an appropriate prior distribution (see, e.g., West and Harrison, 1997; Prado et al., 2021).
@@
The model in (1) delivers inference through MCMC or forward filtering-backward sampling (Carter and Kohn, 1994; Frühwirth-Schnatter, 1994). However, with high-dimensional parameters, these methods require substantial computational resources that render their practical application to be challenging or even infeasible. Consequently, we devise Bayesian predictive stacking that exploits analytically tractable posterior distributions.
@@ 2.2. Dynamic trajectory model
$y_{t}(\gamma(t))={x}_{t}(\gamma(t))^{\mathrm{\scriptscriptstyle T}}{\beta}_{t}+w_{t}(\gamma(t))+\eta_{1t}(\gamma(t)),\quad\eta_{1t}(\gamma(t))\overset{\mathrm{i.i.d.}}{\sim}N(0,\sigma^{2})\;,$ (2)
where ${x}_{t}(\gamma(t))$ is a $p$-dimensional explanatory variable, ${\beta}_{t}$ is a $p$-dimensional time-varying regression coefficient, $w_{t}(\gamma(t))$ and $\eta_{1t}$ are zero-centered spatial and white noise processes, respectively, at time $t$, and $\sigma^{2}$ is the variance of the white noise process. The domain of the process $w_{t}(\cdot)$ is not Euclidean, but an arbitrary trajectory defined by a string of coordinates mapped by $\gamma(t)$. Furthermore, a subject may revisit the same location multiple times, yielding multiple values at $\gamma(t)$, making the spatial covariance matrix singular and, therefore, precluding legitimate probabilistic inference. We obviate this as follows.
Let $\Gamma=\{\gamma(1),\ldots,\gamma(T)\}$ be a complete enumeration of spatial locations visited by the subject, of which $n\leq T$ are distinct spatial locations, and let $\widetilde{\Gamma}=\{\widetilde{\gamma}_{1},\ldots,\widetilde{\gamma}_{n}\}\subseteq\Gamma$ be the subset of distinct locations. Let $z_{t}(\gamma(t))$ be a latent process and map $w_{t}(\gamma(t))=\sum_{j=1}^{n}b(\gamma(t),\widetilde{\gamma}_{j})z_{t}(\widetilde{\gamma}_{j})$, where $b(\gamma(t),\widetilde{\gamma}_{j}):\Gamma\times\widetilde{\Gamma}\to\{0,1\}$ such that $b(\gamma(t),\widetilde{\gamma}_{j})=1$ if $\gamma(t)=\widetilde{\gamma}_{j}$ and 00 otherwise. This yields $w=Bz$, where $w=(w_{1}(\gamma(1)),\ldots,w_{T}(\gamma(T)))^{\mathrm{\scriptscriptstyle T}}$ is $T\times 1$, ${z}=({z}_{1}^{\mathrm{\scriptscriptstyle T}},\ldots,{z}_{T}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ is $nT\times 1$ with each ${z}_{t}=({z}_{t}(\widetilde{\gamma}_{1}),\ldots,{z}_{t}(\widetilde{\gamma}_{n}))^{\mathrm{\scriptscriptstyle T}}$ being $n\times 1$, and $B$ is $T\times nT$, whose $(t,n(t-1)+j)$th entry is $b(\gamma(t),\widetilde{\gamma}_{j})$. This maps the latent spatial effects, $z$, to those at the observed points, $w$. If ${\beta}=({\beta}_{1}^{\mathrm{\scriptscriptstyle T}},\ldots,{\beta}_{T}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$, then temporal autoregressive models for $\beta$ and $z$ are specified as ${\beta}=(A\otimes I_{p}){\beta}+\eta_{2t}$ and $z=(A\otimes I_{n})z+\eta_{3t}$, respectively, where $\eta_{2t}\overset{\mathrm{i.i.d.}}{\sim}N({0},\sigma^{2}\delta_{\beta}^{2}(I_{T}\otimes W_{p}))$ and $\eta_{3t}\sim N(0,\sigma^{2}\delta_{z}^{2}(I_{T}\otimes K_{\phi}))$, $A=\left(\begin{smallmatrix}{0}^{\mathrm{\scriptscriptstyle T}}&0\\ I_{T-1}&0\\ \end{smallmatrix}\right)\in\mathbb{R}^{T\times T}$, $W_{p}\in\mathbb{R}^{p\times p}$ is a correlation matrix among the coefficients and $\otimes$ denotes the Kronecker product. We construct the following augmented model,
@@
3. Continuous space-time trajectory model
We can treat actigraph data as a partial realization of a continuous spatial-temporal process. We write $y(\gamma(t),t)$ to be the measurement that can exist, conceptually, at any time $t\in\mathbb{R}^{+}$ and spatial location $\gamma(t)\in\mathbb{R}^{2}$ at $t$. We construct a regression model.
@@ 3. Continuous space-time trajectory model
over a finite set of $n$ space-time coordinates $(\gamma(t),t)$, where ${x}(\gamma(t),t)$ is $p\times 1$ consisting of explanatory variables, ${\beta}(t)$ is the corresponding $p\times 1$ vector of slopes, $z(\gamma(t),t)$ is a zero-centered spatial-temporal process and $\eta_{1}(t)$ is the measurement error distributed as a zero-centered Gaussian distribution with variance $\sigma^{2}$.
@@
where $K_{\phi}$ is a valid spatial-temporal correlation kernel (Cressie and Huang, 1999; Gneiting, 2002; Stein, 2005). Among rich classes of functions, we work with a non-separable function,
@@
which is a special case of a more general class in Gneiting (2002). For each regression coefficient, we consider the following process:
@@
with temporal correlation kernel $C_{\xi}(t,t^{\prime})=\exp(-\xi^{2}|t-t^{\prime}|^{2})$. We note that $K_{\phi}$ is a positive-definite kernel ensuring that the stochastic process (5) is well-defined. In particular, it is crucial to note that even when $\gamma(t)=\gamma(t^{\prime})$, i.e., the subject returns to the same location at a later time, the function $(1+\phi_{1}|t-t^{\prime}|^{2})^{-1}$ is positive-definite. This is seen from
@@
Since $|t-t^{\prime}|^{2}$ and $1$ are conditionally negative-definite, the integrand is positive-definite from Schoenberg’s theorem (e.g., Phillips et al., 2019), which implies that $(1+\phi_{1}|t-t^{\prime}|^{2})^{-1}$ is also positive-definite. If we observe $y$ at space-time points $(\Gamma,\mathcal{T})=\{(\gamma(t_{i}),t_{i})\mid i=1,\ldots,n\}\subset\mathbb{R}^{2}\times\mathbb{R}^{+}$, then (4),(5) and (7) for $n$ space-time points yields:
@@
where each ${x}_{j}\in\mathbb{R}^{n\times n}$ is diagonal with entries $x_{j}(\gamma(t_{i}),(t_{i}))$ for $i=1,\ldots,n$ and $j=1,\ldots,p$; $X$ is $3n\times(p+1)n$, ${\theta}$ is $(p+1)n\times 1$ consisting of the $pn\times 1$ vector of regression coefficients $\beta=(\beta_{1}^{\mathrm{\scriptscriptstyle T}},\ldots,\beta_{p}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ and the $n\times 1$ vector $z=(z(\gamma(t_{1}),t_{1}),\ldots,z(\gamma(t_{n}),t_{n}))^{\mathrm{\scriptscriptstyle T}}$, $K_{\phi}(\Gamma,\mathcal{T})=(K_{\phi}((\gamma(t_{i}),t_{i}),(\gamma(t_{j}),t_{j})))$ and $C_{\xi}(\mathcal{T})=(C_{\xi}(t_{i},t_{j}))$ are both $n\times n$, where $i,j=1,\ldots,n$ and $S=I_{n}\oplus\left(I_{p}\otimes\delta_{\beta}^{2}C_{\xi}(\mathcal{T})\right)\oplus\delta_{z}^{2}K_{\phi}(\Gamma,\mathcal{T})$. We further assign the prior distribution $\sigma^{2}\sim IG(a_{\sigma},b_{\sigma})$, where $a_{\sigma}$ and $b_{\sigma}$ are fixed rate and scale parameters for the inverse-Gamma distribution. As in (3), the prior distribution for $\theta$ is absorbed into (8) and the posterior distribution for $\{\theta,\sigma^{2}\}$ for any fixed set $\{\delta_{\beta},\delta_{z},\xi,\phi\}$ is in the Normal-IG family. We exploit these distributions to devise stacked inference for $\beta_{j}(t)$ and $z(\gamma(t),t)$.